you are given a polynomial equation and one or more of its roots. Find the remaining roots.
19. 2x^3 - 5x^2 - 4x + 3=0 ; root : x=3
please explain solutions
I would really appreciate any help
please help soon

Answer by stanbon(75887)

2x^3 - 5x^2 - 4x + 3 = 0 ; root : x=3
--------------------------
Use synthetic division to divide by x-3
3)....2....-5....-4....3
.......2.....1.....-1..|.0
Quotient: 2x^2 + x -1 = 0
Factor: (2x-1)(x+1) = 0
x = 1/2, or x = -1 are the other roots
=======================================
Cheers,
Stan H.